# Is the panel fair? Evaluating panel compositions through network analysis. The case of research assessments in Italy

### v2

@@ Abstract
In research evaluation, the fair representation of panels is usually defined in terms of observable characteristics of scholars such as gender or affiliations. An empirical strategy is proposed for exploring hidden connections between panellists such that, despite the respect of formal requirements, the panel could be considered alike as unfair with respect to the representation of diversity of research approaches and methodologies. The case study regards the three panels selected to evaluate research in economics, statistics and business during the Italian research assessment exercises. The first two panels were appointed directly by the governmental agency responsible for the evaluation, while the third was randomly selected. Hence the third panel can be considered as a control for evaluating about the fairness of the others. The fair representation is explored by comparing the networks of panellists based on their co-authorship relations, the networks based on journals in which they published and the networks based on their affiliated institutions (universities, research centres and newspapers). The results show that the members of the first two panels had connections much higher than the members of the control group. Hence the composition of the first two panels should be considered as unfair, as the results of the research assessments.
@@ 1 Introduction
Research evaluation has now become one of main instruments for university and research policies. It is no longer considered as just an essential step for distributing resources to competing research projects and for hiring and promoting researchers; it is also the core of the administrative processes labelled “performance based research funding” where ex-post evaluation is used for financing universities and other research institutions [Langfeldt, 2004, Hicks, 2012, Zacharewicz et al., 2018]. As a consequence, research evaluation has gained centrality in the scientific knowledge-making process, so much that it has taken “the function of gatekeeping, filtering, and legitimating” it [Lamont and Huutoniemi, 2011, p. 209].
Despite their procedural variety, evaluation is usually governed by panels of experts or peers that managed the processes of peer review [Whitley et al., 2010]. Non-scientific characteristics of these panels, like sex, age, or affiliation, may affect evaluation practices, with a concrete risk that specific groups can capture regulatory processes of scientific evaluation systems. This problem has been mainly analyzed in reference to the ex-ante research evaluation, specifically in reference to editorial evaluation adopted by journals and to evaluation of projects for funding [Cole, 1992, Cole and Cole, 1981, Cole et al., 1979, Lamont and Huutoniemi, 2011, Rahman et al., 2016]. Only a few studies have been done on the ex-post research evaluation, which is referred to national research assessments such us the ones applied in United Kingdom and Italy [Lee, 2007, Lee et al., 2013, Harley and Lee, 1997, Baccini, 2016c, b, Corsi et al., 2010, 2011].
The central notion in these discussions is procedural fairness, which should be “concerned with procedures used to arrive at [fair] outcomes” [Beersma and De Dreu, 2003, p. 220] and that includes the fair representation of all affected parties involved in the decision-making process [Leventhal, 1980]. The concrete applications of this procedural fairness is present in the European Peer Review Guide of the European Science Foundation [ESF, 2011] and it was adopted in the composition of panels of British Research Excellence Framework (REF) and also of the Italian research assessments. However, the problem is how to define who the affected parties are and what dimensions to take into account in defining their fair representation. The usual way is to define observable characteristics of scholars, such as research field, gender, affiliation and geography. But if a discipline is characterized by the coexistence of many schools of thought with different approaches, methodologies and policy recipes, a fair composition in terms of observable characteristics may not guarantee a fair representation of the different schools of thought. The unfair selection of panellists may contribute to the prevalence of one of competing paradigms or schools of thought, by reinforcing the normative standard, pre-existing journal rankings and more generally the hierarchy of a discipline.
@@
Probably the most studied case is economics (see Corsi et al. [2019] for a review of the literature), also because the evaluation of economic theories affects economic policy [Ferguson and Johnson, 2018]. Pioneer of these studies was Frederic S. Lee, who analyzed and harshly criticised the Research Assessments Exercise (RAE) in the United Kingdom [Harley and Lee, 1997, Lee, 2007, Lee et al., 2013]. He explicitly individuated a precise role for the panel of RAE, composed by paradigmatic homogeneous experts “controlled by mainstream economists [who] have used [the RAE] to support particular neoclassical research over heterodox research and promote neoclassical departments over more pluralistic ones” [Lee, 2007, p.15]. A process aimed “to achieve a discipline-desired outcome that was (and is) compatible with the Government’s pro-market ideological agenda” [Lee, 2007, p.14].
The Italian case also has been considered as an internationally relevant example on how it can disregard heterodox schools and historical methods in favour of mainstream approaches and quantitative methods in economics [Pasinetti, 2006, Corsi et al., 2010, 2011]. Re [2019] claimed that this process was connected to a cultural and political change from the Keynesian to the Ordoliberal ideology. The lack of fairness in the composition of the committees in charge of economic evaluation was also highlighted: their homogeneity could probably have minimized the voices of disagreement with the evaluation methods and rules adopted [Pasinetti, 2006, Baccini, 2011, 2014, 2016c, Baccini and Ricciardi, 2012].
@@
This paper proposes an empirical strategy for exploring if, despite formal adherence to fair panel composition in terms of easily observable characteristics, it is possible to detect connections among members such that panel composition can be considered as unfair. The case studies regards the composition of the panels appointed to evaluate research in economics, statistics and business during the Italian research assessment exercises, the so called VQR 2004-2010, VQR 2011-2014 and VQR 2015-2019. The Italian case is of particular interest since the three considered panels were selected by ANVUR, the governmental agency for the evaluation of university and research, by using different rules. Indeed, as it is detailed later on, while the first two panels were appointed directly by the governing board of ANVUR, the third was instead selected randomly by a lot. Hence the third panel can be considered as a sort of control group [Gillies, 2014]. For investigating the fairness of panel composition, the social and intellectual connections between panel members are investigated through network analysis techniques. Three kinds of connections, and hence three kinds of networks are considered and explored: (i) the coauthorship networks built by considering the publications of the members of the panels; (ii) the journal based networks built by considering the journals where the panellists publish their scholarly articles; and (iii) the “affinity networks” built by considering panellists’ affiliations to universities and research centres, and their collaborations to newspapers or blogs.
@@
The organization of the article is as follows: Section 2 reviews literature on procedural fairness; in Section 3, the Italian research assessments and the panel selection methods are described; in Section 4 the research design is presented; sections 5-7 report the analysis, respectively, of the co-authorship networks, of the journal based networks, and of the affinity networks. Section 8 discusses results and concludes.
2 The notion of fair representation in research evaluation
@@
It has been shown that some of the perverse effects of peer review, such as cronyism, the pursuit of self-interest, and cognitive particularism, may be influenced by the way panels are set up [Lamont and Huutoniemi, 2011]. Both potential and observed risks in the peer review system have been discussed and documented. It is argued that the system is conservative and suppresses innovative research. Several scholars have even suggested that peer review hinders scientific progress [Horrobin, 1990]. Effects such as nepotism and old-boyism in peer review are seen to hinder pioneering research [Chubin and Hackett, 1990, Roy, 1985], while “cognitive particularism”, “favoritism for the familiar” and “scholarly bias” support the school viewpoint or research topic the reviewers themselves are conducting (see, e.g., Porter and Rossini [1985], Travis and Collins [1991], Langfeldt [2004]). The relations between panellists and evaluated scholars is a channel of bias: Wennerås and Wold [1997] found that higher competence scores are given to applicants who are affiliated with a panellist than to applicants without such ties. Moreover, another possible group effect is groupthink, which refers to “a deterioration of mental efficiency, reality testing, and moral judgement that results from in-group pressures” [Janis, 1982, p.9]. Loyalty to the group “requires each member to avoid raising controversial issues, questioning weak arguments, or calling a halt to soft-headed thinking” [Janis, 1982, p.12]. Wennerås and Wold [1997] also argued that peer review fosters nepotism and gender bias, disadvantaging women, though subsequent studies have not consistently supported these claims Marsh et al. [2009]. However, other studies suggest that gender dynamics in academic evaluations may indeed play a role: Van den Brink [2010] found that in the Netherlands, an increased presence of women on appointment committees led to more women being appointed as full professors, suggesting a preference for same-gender candidates.
As a solution to these problems, some studies [Roy, 1985, Bell, 1992, GAO, 1994] focused on the fairness of the peer review process showing that fair process of evaluating knowledge can be put in place once particularistic considerations are eliminated. Lamont and Huutoniemi [2011] claimed instead that extra-cognitive factors do not corrupt the evaluation process but are intrinsic to it and that the “fairness of the process is not undermined by non-rational features but is created through inter-subjective rules that evaluators follow to distinguish between legitimate and illegitimate behavior”. For this reason, studies on evaluation and peer reviews focused on the procedural fairness, which “is concerned with procedures used to arrive at those outcomes” [Beersma and De Dreu, 2003, p.220] and not solely on the allocative results. The concept of distributive fairness restricts the analysis of perceived justice only to the final distribution of rewards or punishments. In the procedural fairness people perceive fairness not solely in terms of the distribution of reward but also in terms of the social system which generates that distribution. “For example, fairness may be judged in terms of a procedure’s consistency over time and across persons; its accuracy and prevention of personal bias; or its representativeness of the values, interests, and outlook of important subgroups in the population of persons affected by the allocative process” [Leventhal, 1980, p.54]. Thus, among these factors there is also the fair representation of all affected parties involved in the decision-making process. In the case of research assessment, this means to focus on the fairness of the composition of panels called to evaluate research.
@@
In fact the role of panellists is very similar to that of the members of a popular jury in a trial [Baccini and Ricciardi, 2012]. In order to have a fair judgment by a panel of judges, it is necessary to designate a fair jury and therefore presumably less inclined to partiality.
Jury selection is the legal procedure in which randomization is most widely used [Duxbury, 2002]. Its use is justified as a way to give a fair chance to every qualified citizen to serve on a jury, and to provide defendants and litigants to be tried by a “representative cross-section of the population” [Duxbury, 2002, p. 75]. The rules of the United States Jury Selection Service Act, for example, state that the jury must be appointed by selecting “at random from a fair cross-section of the community”. The fair composition of the popular jury aims to ensure fairness of the judgment. “Achieving representative cross-sections of the community in jury venires, and ensuring that our civil juries reflect the community as well, are essential components contributing to the fairness and legitimacy of our civil justice system [..] and the representativeness of juries is not merely an aspiration but a guarantee under state and federal constitutions and statutes” [Hans, 2021, p.1].
@@
In particular, it has been shown that juries that reflect the full range of community perspectives are in a position to incorporate these diverse views into their fact finding. The best-known and best-documented examples concern the need to balance popular juries from the point of view of ethnic groups because it is believed that a jury composed mainly of members of the same ethnic group tends to be favourable towards a defendant of the same group, and unfavourable to a defendant from a different ethnic group. For example, Sommers and Ellsworth [2003] in a mock jury experiment, comparing the deliberations of all-white and racially mixed juries, discovered that diverse jury deliberations were more accurate, more expansive, and longer. It was not simply that the minority jurors contributed new and different information, the white jurors acted differently in all-white versus mixed-race juries: they made fewer factual mistakes, and raised more issues and evidence, during the deliberation. Moreover, representative juries are more likely to be seen as legitimate decision makers, which in turn contributes to public confidence in the justice system. For all these reasons, “courts should ensure that jury selection procedures serve the goal of maximizing the representativeness of jury pools and civil juries” [Hans, 2021, p.8].
@@
Similar considerations can be found in documents related to the design of research assessments. For instance, in the document regarding the recruitment of panels for the British RAE, it is recommended a fair representation of “all affected parties” involved in the decision-making process. It was stated that the selection of panellist has to ensure that “the overall body of members reflects the diversity of the research community, including in terms of age, gender, ethnic origin, scope and focus of their home institution, and geographical location which represents the international reference on the subject” [REF, 2010]. The European Peer Review Guide of the European Science Foundation suggests that “the goal should be to ensure availability of diverse viewpoints, scientific perspectives and scholarly thinking” and that the criteria to be adopted for the selection of experts must also be the ‘diversity’ that is expressed in terms of “gender balance, scholarly thinking, background, geography, turnover” [ESF, 2011].
The basic tenet of this kind of reasoning is that, as for legal juries, a panel composed by a cross-section of members of a scientific community may reach not only procedurally fair decisions, but also substantively better decisions [Elster, 1989, p. 97].
@@
Indeed, the fairness of the composition of panels for the VQR 2004-2010 was questioned from its inception, by highlighting the lack of transparency in the members appointing procedures [Baccini, 2011]. The attention was especially focused on the panel for economics, statistics and business. This panel was composed not only by a small minority of women (16.7%), but its appointed members were closely linked to each other by co-authorship relationships. In the economics sub-panel, 9 panellists out of 20 (45%) were among the founders of a ultra-liberal party (Fare per fermare il declino) that participated in the political elections of 2013 by obtaining the 0.9% of the votes and no representatives in parliament [Baccini, 2018]. The same problem was also documented for the VQR 2011-2014 during a conference organised at the Italian Parliament: the panel of economics was again largely composed by scholars of the same ultra-liberal party [Baccini, 2016a]. In the public debate it was noted that the absence of fairness in the composition of the panel for economics, statistics and business had probably minimized the voices of dissent with respect to the evaluation methods and rules adopted by the panel [Baccini, 2011, Re, 2019]. It was noted also that there was a relevant precedent to consider: in the first experimental research assessment for the years 2001-2003, managed by CIVR (Steering Committee for Research Evaluation), Luigi Pasinetti, one of the panel member for economics, wrote a note of dissent documenting the absence of pluralism in evaluation [Pasinetti, 2006].

### v3

@@ Abstract
Research evaluation is usually governed by panels of peers. Procedural fairness refers to the principles that ensures decisions are made through a fair and transparent process. It requires that the composition of panels is fair. A fair panel is usually defined in terms of observable characteristics of scholars such as gender or affiliations. The formal adherence to these criteria is not sufficient to guarantee a fair composition in terms of scholarly thinking, background, or policy orientation. An empirical strategy for exploring the fairness in the intellectual composition of panels is proposed, based on the observation of links between panellists. The case study regards the three panels selected to evaluate research in economics, statistics and business during the Italian research assessment exercises. The first two panels were appointed directly by the governmental agency responsible for the evaluation, while the third was randomly selected. Hence the third panel can be considered as a control for evaluating about the fairness of the others. The fair representation is explored by comparing the networks of panellists based on their co-authorship relations, the networks based on journals in which they published and the networks based on their affiliated institutions (universities, research centres and newspapers). The results show that the members of the first two panels had connections much higher than the members of the control group. Hence the composition of the first two panels should be considered as unfair, as the results of the research assessments.
@@ 1 Introduction
Research evaluation has now become one of main instruments for university and research policies. It has been “integrated as administrative routine at many levels and in many types of institutions” [Dahler-Larsen, 2011, p. 2]. It is no longer considered as just an essential step for distributing resources to competing research projects and for hiring and promoting researchers; it is also the core of the administrative processes labelled “performance based research funding” where ex-post evaluation is used for financing universities and other research institutions [Langfeldt, 2004, Hicks, 2012, Zacharewicz et al., 2018]. The consequences are twofold. First, research evaluation has gained centrality in the scientific knowledge-making process, so much that it has taken “the function of gatekeeping, filtering, and legitimating” it [Lamont and Huutoniemi, 2011, p. 209]. Second, research evaluation, being entered in administrative processes, should follow their rules, and among these the ones regarding procedural fairness [Ruder and Woods, 2019].
Procedural fairness is “concerned with procedures used to arrive at [fair] outcomes” [Beersma and De Dreu, 2003, p. 220], i.e. procedures should be “designed to reduce bias or favoritism in government decision making” [Ruder and Woods, 2019, p. 400]. In general, procedural fairness aims to protect individuals’ rights and maintain trust in the integrity and fairness of administrative decision-making. Procedural fairness in an administrative procedure refers to the principles that ensures decisions are made through a fair and transparent process. It typically involves the right to an unbiased decision-maker: decisions must be made by an impartial authority who does not have any personal interest, bias, or preconceptions that could influence the outcome [Leventhal, 1980, Bobocel and Gosse, 2015].
Despite their concrete variety, research evaluation is usually governed by panels of experts or peers that managed or directly realized the processes of peer review [Whitley et al., 2010]. Hence, the composition of these panels is a key issue for procedural fairness, which generally requires fair representation of all affected parties involved in the decision-making process [Leventhal, 1980]. The problem is how to define who the affected parties are and what dimensions to take into account in defining their fair representation. The usual way in which this problem is handled is by considering observable non-scientific characteristics of scholars, such as gender, age, affiliation, geography that may influence evaluation practices. Also other scientific characteristics of scholars are considered, such as their field of research or scientific prominence in order to ensure their ability to make an expert evaluation. The building of fair panels requires a proper balance of all these characteristics. The European Peer Review Guide of the European Science Foundation [ESF, 2011], the rules for the composition of panels of British Research Excellence Framework (REF) and of the Italian research assessments are examples of the way in which institutions concretely try to build fair evaluation panels.
An unbalanced panel composition is usually seen as a real risk that certain groups will capture the evaluation systems. Many studies have examined how panel composition affects evaluation outcomes and the resulting distribution of awards. They analyzed evaluation procedures in view of showing the respect or the departure from the Mertonian universalism [Merton, 1973], according to which the evaluation and the distribution of awards should reflect the distribution of “intellectual deservingness of recipients, rather than to their group identity or status or their particularistic relationship with the panellists” [Mallard et al., 2009, 575]. In doing so, these studies focused on the relationship between the distribution of awards or results of evaluation and the social characteristics (e.g., age, race, gender, scientific prominence) of panellists or of individuals whose work is evaluated. They regarded mainly cases of ex-ante research evaluation [Cole, 1992, Cole and Cole, 1981, Cole et al., 1979, Lamont and Huutoniemi, 2011, Rahman et al., 2016], but also cases of ex-post research evaluation, such as the national research assessments in United Kingdom and Italy [Lee, 2007, Lee et al., 2013, Harley and Lee, 1997, Baccini, 2016c, Baccini and De Nicolao, 2021, Corsi et al., 2010, 2011].
The research about “procedural fairness” in research evaluation has considered also the processes through which panels arrive at final decisions [Mallard et al., 2009]. The specific composition of a panel contributes to the definition of the “set of informal, and sometimes inconsistent, procedural rules that the panellists tacitly co-construct and then adhere to in the course of their deliberations” [Camic et al., 2011, p. 18]. Hence, the way panels are set up may generate cronyism, the pursuit of self-interest, and cognitive particularism in peer review [Lamont and Huutoniemi, 2011]. In this literature, the relevant dimension in the composition of a panel is disciplinary, because scientists are socialized in evaluative practices which are discipline specific. These practices are the basis from which panel members interact to reach evaluative deliberations [Lamont and Huutoniemi, 2011]. In sum, a fair composition of a panel in terms of observable characteristics of its members is not sufficient to guarantee that the composition of the panel is also, so to speak, “intellectually” fair. The lack of a fair intellectual composition can be considered as a way in which procedural unfairness emerges in research evaluation. Similarly, when a panel is called to evaluate a discipline characterized by the coexistence of many schools of thought with different approaches, methodologies, policy recipes and evaluative practices, its sub-disciplinary composition may be relevant to its deliberations. In this case an intellectual unfair selection of panellists may contribute to the prevalence of one of competing paradigms or schools of thought, by reinforcing the normative standard, pre-existing journal rankings and more generally the hierarchy of a discipline.
@@
Probably, economics is the most studied case of the dysfunctional effects [Ferguson and Johnson, 2018] of unfair compositions of disciplinary panels in research evaluation (see Corsi et al. [2019] for a review of the literature). Pioneer of these studies was Frederic S. Lee, who analyzed the Research Assessments Exercise (RAE) in the United Kingdom [Harley and Lee, 1997, Lee, 2007, Lee et al., 2013]. He explicitly individuated a precise role for the panel of RAE, composed by paradigmatic homogeneous experts “controlled by mainstream economists [who] have used [the RAE] to support particular neoclassical research over heterodox research and promote neoclassical departments over more pluralistic ones” [Lee, 2007, p.15]. A process aimed “to achieve a discipline-desired outcome that was (and is) compatible with the Government’s pro-market ideological agenda” [Lee, 2007, p.14]. The Italian case also has been considered as an internationally relevant example on how it can disregard heterodox schools and historical methods in favour of mainstream approaches and quantitative methods in economics [Pasinetti, 2006, Corsi et al., 2010, 2011], probably connected to a cultural and political change from the Keynesian to the Ordoliberal ideology [Re, 2019]. The lack of fairness in the composition of the committees in charge of economic evaluation was also highlighted: their homogeneity could probably have minimized the voices of disagreement with the evaluation methods and rules adopted [Pasinetti, 2006, Baccini, 2011, 2014, 2016c, Baccini and Ricciardi, 2012].
@@
This paper proposes an empirical strategy for reasoning about the fairness in the intellectual composition of panels. It attempts to detect connections among members such that panel composition can be considered as unfair, despite formal adherence to a fair panel composition in terms of easily observable characteristics such as gender, affiliation or geography. The social and intellectual connections between panel members are investigated through network analysis techniques. Three kinds of connections, and hence three kinds of networks are considered and explored: (i) the coauthorship networks built by considering the publications of the members of the panels; (ii) the journal based networks built by considering the journals where the panellists publish their scholarly articles; and (iii) the “affinity networks” built by considering panellists’ affiliations to universities and research centres, and their collaborations to newspapers or blogs.
The main difficulty of this kind of approach is to define a threshold of connections above which the panel composition should be considered unfair. The basic intuition is that the panel should be considered unfair if the connections between its members are anomalously higher than the normal connections that occur in the population of scholars from which the panel was chosen. Unfortunately, this threshold can be difficult, if not impossible, to calculate in practice, if the population of scholars is too large or not clearly defined. The case study developed here regards the composition of the panels appointed to evaluate research in economics, statistics and business during the Italian research assessment exercises, the so called VQR 2004-2010, VQR 2011-2014 and VQR 2015-2019. The three panels were selected by ANVUR, the governmental agency for the evaluation of university and research, by using different rules. Indeed, as it is detailed later on, the first two panels were appointed directly by the governing board of ANVUR; the third panel was instead selected randomly by a lot. Hence this third panel can be considered as a sort of control group: the connections between members of the third panel can be considered as the reference threshold for judging the fairness of the composition of the other two panels.
@@
The organization of the article is as follows: Section 2 reviews literature on procedural fairness and panel composition; in Section 3, the Italian research assessments and the panel selection methods are described; in Section 4 the research design is presented; sections 5-7 report the analysis, respectively, of the co-authorship networks, of the journal based networks, and of the affinity networks. Section 8 discusses results and concludes.
2 Research evaluation, procedural fairness and panel composition
@@
In a literature review, Mallard et al. [2009] stated that studies of peer review and research evaluation had until then largely focused on distributive fairness by restricting the analysis only to the final results of evaluation and the consequent distribution of rewards or punishments. Indeed, the so called “first wave” of research on peer review and research evaluation [Lamont and Mallard, 2005] was mainly focused on obstacles to “fairness caused by nonscientific influences such as politics, friendship networks, or common institutional positions” [Travis and Collins, 1991]. Both potential and observed risks in the peer review system were discussed and documented. It was argued that the system is conservative and suppresses innovative research. Several scholars even suggested that peer review hinders scientific progress [Horrobin, 1990]. Effects such as nepotism and old-boyism in peer review were seen to hinder pioneering research [Chubin and Hackett, 1990, Roy, 1985], while “cognitive particularism”, “favoritism for the familiar” and “scholarly bias” support the school viewpoint or research topic the reviewers themselves are conducting (see, e.g., Porter and Rossini [1985], Travis and Collins [1991]). Moreover, another possible group effect is groupthink, which refers to “a deterioration of mental efficiency, reality testing, and moral judgement that results from in-group pressures” [Janis, 1982, p.9]. Loyalty to the group “requires each member to avoid raising controversial issues, questioning weak arguments, or calling a halt to soft-headed thinking” [Janis, 1982, p.12].
The empirical literature of this first wave of research largely that universalistic norms were followed more often than not [GAO, 1994, Cole et al., 1979, Cole and Cole, 1981, Merton and Zuckerman, 1971]. Subsequent studies produced mixed evidence. Wennerås and Wold [1997] argued that peer review fosters nepotism and gender bias, disadvantaging women, though subsequent studies have not consistently supported these claims [Marsh et al., 2009]. Other studies suggest that gender dynamics in academic evaluations may indeed play a role: Van den Brink [2010] found that in the Netherlands, an increased presence of women on appointment committees led to more women being appointed as full professors, suggesting a preference for same-gender candidates. Langfeldt [2004] documented a variety of source of bias in panels appointed to evaluate research. These biases emerged from interactions (discussion, negotiations, lack of agreement) between panel members, for which the selection of panel was crucial.
According to Mallard et al. [2009], the first wave of studies did not systematically consider the question of “procedural fairness”, i.e. the fairness of the procedures used in research evaluation to produce evaluative results. That question started to be explored by a second wave of literature on research evaluation and peer review, focusing in particular on cognitive dimensions of peer evaluation. Collins and Evans [2002] discussed the nature and role of scientific expertise in decision making processes. Lamont and Mallard [2005], Mallard et al. [2009], Lamont and Huutoniemi [2011] proposed a very specific interpretation of procedural fairness, as the way in which panels of evaluators concretely construct the evaluation process, i.e. how they define and apply rules and criteria of evaluation. Lamont and Huutoniemi [2011] analyzed the customary rules constructed by some panellists in research evaluation and how they perceive the outcome of their decisions as “fair”. Panellists tend to perceive their decision as fair in reference to the concrete rules and process they used for handling disagreement and negotiating agreement. panellists consider fair an evaluative procedure if they shared the “belief that meritocracy guides the process, while corrupting forces, self-interest, and, in particular, politics are kept at bay” [Lamont and Huutoniemi, 2011, p. 214]. In summary, according to this literature, research evaluation is shaped and constrained by the procedural context in which occurs. It is therefore useless to contrast biased and unbiased evaluation, because “extra-cognitive factors do not corrupt the evaluation process but are intrinsic to it” [Lamont and Huutoniemi, 2011, p. 228]. In this regard, the structure and composition of panels is of primary interest because it can influence the “customary rules” concretely adopted in research evaluation. An “unfair evaluation” may result from a panel composition in which evaluators agree to adopt rules that favor research similar to their ones in terms of methodological or epistemological features, research styles or policy prescriptions.
The focus on procedural fairness can be understood in reference also to different streams of literature regarding the institutional frame of research evaluation and peer review. The first step of the reasoning consists in considering that research evaluation and peer review are integrated as administrative routine in many types of institutions [Dahler-Larsen, 2011]. Indeed, research evaluation is carried out by governments, academic institutions, agencies, or other organizations to produce indicators or other evaluation information. Pievatolo [2024, p. 6] highlighted that in the administrative evaluation of research, “assessment authorities” give to evaluators decision-making power. When scholars are in charge of evaluation in an administrative procedure, they are subtracted to the public scrutiny of their opinion as usually happen in the scientific debate, because they can impose their decisions “even on those who disagree”[Pievatolo, 2024, p. 6]. Hence, if research evaluation has fully entered into administrative activities such as recruitment, selection, funding, it will eventually follow its rules. Procedural fairness - sometimes also referred to as procedural justice or procedural due process - is a cornerstone of administrative processes. They are fair when they are designed to reduce bias or favoritism in decision making. Fairness may be judged in terms of a procedure’s “consistency over time and across persons; its accuracy and prevention of personal bias; or its representativeness of the values, interests, and outlook of important subgroups in the population of persons affected” [Leventhal, 1980, p.54]. Transparency and public participation are elements of procedural fairness. It is the main source of legitimacy of administrative decisions made by government agency [Ruder and Woods, 2019]. It is a signal that people subjected to the decision are respected, that decision-makers can be trusted and that the final decision can be considered fair [Esaiasson et al., 2019, p. 293]. According to Frey et al. [2004], procedural fairness generates procedural utility because rational people have preferences about how outcomes are generated.
@@
If we consider research evaluation as a special kind of administrative procedure, its fairness concerns also the composition of evaluation panels. Baccini and Ricciardi [2012] suggested an analogy between the role of panellists in research evaluation and that of members of a popular jury in a trial. In order to have a fair legal judgment by a panel of judges, it is necessary to designate a fair jury and therefore presumably less inclined to partiality. Similarly, in order to have a fair research evaluation, the panel should be composed in a fair manner.
@@
The fair procedures to select a jury is a widely debated issue. Randomization is considered as an effective way to build a fair jury, by giving a fair chance to every qualified citizen to serve on a jury, and to provide defendants and litigants to be tried by a “representative cross-section of the population” [Duxbury, 2002, p. 75]. The rules of the United States Jury Selection Service Act, for example, state that the jury must be appointed by selecting “at random from a fair cross-section of the community”. The fair composition of the popular jury aims to ensure fairness of the judgment. “Achieving representative cross-sections of the community in jury venires, and ensuring that our civil juries reflect the community as well, are essential components contributing to the fairness and legitimacy of our civil justice system [..] and the representativeness of juries is not merely an aspiration but a guarantee under state and federal constitutions and statutes” [Hans, 2021, p.1]. In particular, it has been shown that juries that reflect the full range of community perspectives are in a position to incorporate these diverse views into their fact finding. The best-known and best-documented examples concern the need to balance popular juries from the point of view of ethnic groups because it is believed that a jury composed mainly of members of the same ethnic group tends to be favourable towards a defendant of the same group, and unfavourable to a defendant from a different ethnic group. For example, Sommers and Ellsworth [2003] in a mock jury experiment, comparing the deliberations of all-white and racially mixed juries, discovered that diverse jury deliberations were more accurate, more expansive, and longer. It was not simply that the minority jurors contributed new and different information, the white jurors acted differently in all-white versus mixed-race juries: they made fewer factual mistakes, and raised more issues and evidence, during the deliberation. Moreover, representative juries are more likely to be seen as legitimate decision makers, which in turn contributes to public confidence in the justice system. For all these reasons, “courts should ensure that jury selection procedures serve the goal of maximizing the representativeness of jury pools and civil juries” [Hans, 2021, p.8].
@@
Similar considerations can be found in documents related to the design of research assessments. For instance, in the document regarding the recruitment of panels for the British RAE, it is recommended a fair representation of “all affected parties” involved in the decision-making process. It was stated that the selection of panellist has to ensure that “the overall body of members reflects the diversity of the research community, including in terms of age, gender, ethnic origin, scope and focus of their home institution, and geographical location which represents the international reference on the subject” [REF, 2010]. The European Peer Review Guide of the European Science Foundation suggests that “the goal should be to ensure availability of diverse viewpoints, scientific perspectives and scholarly thinking” and that the criteria to be adopted for the selection of experts must also be the ‘diversity’ that is expressed in terms of “gender balance, scholarly thinking, background, geography, turnover” [ESF, 2011]. It appears that, as for legal juries, a panel composed by a cross-section of members of a scientific community may reach not only procedurally fair decisions, but also substantively better decisions [Elster, 1989, p. 97]. From an empirical point of view, it is easy to test the fairness of a panel’s composition in terms of observable characteristics of its members, such as gender, age or affiliation. It is instead much more complicated to test the fairness of its composition in terms of scholarly thinking, background, or other relevant features. In what follows, a test is proposed for fairness of a panel’s composition based on the observation of intellectual and social connections among its members.
@@
Indeed, the fairness of the composition of panels for the VQR 2004-2010 was questioned from its inception, by highlighting the lack of transparency in the members appointing procedures [Baccini, 2011]. The attention was especially focused on the panel for economics, statistics and business. This panel was composed not only by a small minority of women (16.7%), but its appointed members were closely linked to each other by co-authorship relationships. In the economics sub-panel, 9 panellists out of 20 (45%) were among the founders of a ultra-liberal party (“Fare per fermare il declino”) that participated in the political elections of 2013 by obtaining the 0.9% of the votes and no representatives in parliament [Baccini, 2018]. The same problem was also documented for the VQR 2011-2014 during a conference organised at the Italian Parliament: the panel of economics was again largely composed by scholars of the same ultra-liberal party [Baccini, 2016a]. In the public debate it was noted that the absence of fairness in the composition of the panel for economics, statistics and business had probably minimized the voices of dissent with respect to the evaluation methods and rules adopted by the panel [Baccini, 2011, Re, 2019]. It was noted also that there was a relevant precedent to consider: in the first experimental research assessment for the years 2001-2003, managed by CIVR (Steering Committee for Research Evaluation), Luigi Pasinetti, one of the panel member for economics, wrote a note of dissent documenting the absence of pluralism in evaluation [Pasinetti, 2006].